OpenStudy (anonymous):

Find the equation of the line through (-1, 0) that is parallel to the line 3x + 5y = -1

OpenStudy (anonymous):

hint rearrange the equation 3x +5y = -1 o the form y = mx +c where m is the gradient, can you do that?

OpenStudy (anonymous):

3x+5y=-1 substract 3x both sides---->5y=-1-3x Divide by 5 both sides----->y=(-3/5)x-1/5

OpenStudy (anonymous):

But we haven't finished yet, right? We will have to use this formula: \[y-y_1=m(x-x_1)\] Remember it?

OpenStudy (anonymous):

yes you have to

OpenStudy (anonymous):

yes i remember